OpenStudy (anonymous):

please help i got to help my little sister with her math and i have no idea how o do this!! solve the quadratic equation (2x-2)(x+7)=0

OpenStudy (thefurball):

ok, so since when they are multiplied they equal zero, then either one of them is zero or they both are zero. that yeilds \[2x-2 = 0\] and\[x+7 = 0\] let me know if you need help with those equations.

OpenStudy (thefurball):

Note: you will have two different answers

OpenStudy (anonymous):

how do they equal 0?

OpenStudy (thefurball):

it's like this:\[a * b = 0\] either a or b can be equal to zero, since anything times 0 is equal to 0

OpenStudy (thefurball):

in this case a is 2x-2 and b is x + 7

OpenStudy (anonymous):

so wouldnt 2x-2= 0x and x+7= 7x?

OpenStudy (anonymous):

and that would give you 0X . 7x and that give is 0 right?

OpenStudy (thefurball):

you can't have 0x, because anything times 0 is 0. there will be two answers, since two values of x can make the expression equal to 0. If x=1, then (2x-2)(x+7) = (0)(8) = 0, so 1 is one of the answers. If x = -7, then (2x-2)(x+7) = (-19)(0) = 0, so -7 is another answer. The way you find both answers is you set each term equal to 0 like i did before. 2x-2 =0 and x+7=0.

OpenStudy (anonymous):

okay i am sorry but i dont understand any of that